# Mass and energy cascade in collisionless dark matter flow and relevant constraints on the dark matter particle massA.6

### v3

@@ 5.5 Particle random walk and halo density profile
where $\Gamma(x,y)$ is an upper incomplete gamma function.
Next, we focus on the halo density profile that can be obtained from the particle distribution function $P_{r}$. The well-known core-cusp problem describes the discrepancy between cuspy halo density predicted by cosmological CDM-only simulations and the core density from observations. The predicted halo density exhibits a cuspy inner density $\rho_{r}\propto r^{s}$ with a wide range of slope parameter $s$ between -1.0 and -1.5 (see Introduction 1). There appears to be no consensus on the exact value of the slope $s$ and no solid theory for the density slope $s$. In this paper, we propose that the internal structure of haloes is primarily determined by the radial mass and energy flow (Xu, 2023c, b). In this regard, we have the following considerations:
  1. 1.
    Individual haloes with both non-zero mass flux and non-vanishing energy flow ($\dot{m}_{r}^{net}\neq 0$ in Eq. (41), $\varepsilon\neq 0$). The outward radial flow can be caused by the migration of particles along the radial direction. Inward radial flow can be caused by halo mass accretion. The stable clustering hypothesis requires that the outflow cancels out the inflow such that the net mass flux vanishes (Fig. 9). However, most simulated haloes are non-equilibrium dynamic objects that might have both a nonzero mass flux and non-vanishing energy flow. The density slopes of these haloes depend both on the radial flow and the mass accretion of each halo (Xu, 2023b). This can be the reason for the wide variety of density slopes $s$ for simulated haloes. Such simulated haloes can generally be modeled by a double-$\gamma$ density profile
    $\rho_{r}(r,t)=\rho_{s}(t)\left(\frac{r}{r_{s}}\right)^{\frac{\alpha}{\beta}-2}\exp\left(\frac{1}{\beta}\left(1-{\left(\frac{r}{r_{s}}\right)^{\alpha}}\right)\right),$ (84)
    where $\rho_{s}(t)$ is the density at scale radius $r_{s}(t)$. This four-parameter double-$\gamma$ density profile ($\rho_{s}$, $r_{s}$, $\alpha$, and $\beta$ in Eq. (84)) reduces to the standard three-parameter Einasto profile (($\rho_{s}$, $r_{s}$, and $\alpha$) with $\alpha=2\beta$.
  2. 2.
    Haloes with vanishing net mass flux and non-vanishing energy flow ($\dot{m}_{r}^{net}=0$ and $\varepsilon\neq 0)$). Individual haloes may not satisfy the stable clustering hypothesis exactly and can have random mass flux with $\dot{m}_{r}^{net}\neq 0$. However, in N-body simulations, we can average out the net mass flux by constructing composite haloes from all haloes of the same mass $m_{h}$. This is how we extract the halo density profiles from N-body simulations (Figs. 13 and 14), where relevant quantities are averaged for all haloes with the same mass. For these composite haloes, the net mass flux vanishes, and we only have the energy flow determining the internal structure of the halo such that the inner density follows the -4/3 law. The corresponding double-density profile for haloes of any mass $m_{h}$ (let $\alpha/\beta=2/3$ in Eq. (84) or $\gamma_{1}=2/3$) reads
    where $\beta_{r}\approx 1$ is an amplitude parameter, $\beta$ is a shape parameter, and $r_{s}$ is the scale radius. Here, $\varepsilon_{u}\approx 10^{-7}m^{2}/s^{3}$ is the rate of the energy flow for haloes with characteristic mass $m_{h}^{*}(z)$. While $\varepsilon(m_{h},z)$ is the rate of the energy flow in haloes of any mass $m_{h}$ (Eq. (68)). For small $r$, the inner density reduces to
    $\rho_{r}(r,m_{h},z)=\beta_{r}\varepsilon^{2/3}G^{-1}r^{-4/3}\propto m_{h}^{4/9}a^{-2/3}\quad\textrm{for}\quad r\rightarrow 0.$ (86)
    The redshift dependence is absorbed into the parameter $\varepsilon$. For haloes of characteristic mass $m_{h}^{*}(z)$, there exists a small-scale permanence for the halo density, i.e., the density profiles at different redshifts $z$ converge to a time-unvarying scaling,
    The small-scale permanence is shown and discussed in Fig. 13.
    Figure 14 presents the density profiles from Illustris simulations for (composite) haloes of different masses $m_{h}$ at $z=0$. The solid lines show the average halo density profiles for all haloes with a mass between $10^{\pm 0.1}m_{h}$ from the Illustris simulation, where halo mass $m_{h}$ is between $10^{8}M_{\odot}$ and $10^{14}M_{\odot}$. The thick dashed (straight) line represents the -4/3 scaling laws for haloes of mass $m_{h}$ that involve the rate of cascade $\varepsilon\propto m_{h}^{2/3}$ (Eq. (68)). Here, $\varepsilon=\varepsilon_{u}$ for haloes of characteristic mass $m_{h}^{*}$ (Eq. (68)). The figure demonstrates the -4/3 law that we obtained from the energy cascade in haloes. That scaling was in agreement with the density profiles of haloes of more than six orders of magnitude. The thin dashed lines in both Figs. 14 and 13 present the double-$\lambda$ density profile from Eq. (85) that has good agreement with simulation data.
  3. 3.
    The last scenario is completely virialized haloes with both vanishing net mass flux and vanishing energy flow ($\dot{m}_{r}^{net}=0$ and $\varepsilon=0$). This is the limiting equilibrium state haloes evolve toward but can never reach. The $-4/3$ density slope should still be good in this limiting state with $\varepsilon\rightarrow 0$.
@@ 6 Constraints on dark matter particle mass
Previous sections discuss the mass and energy cascade in halo mass space and in individual haloes, which can be relevant to dark matter particle mass and properties. This section focuses on some potential constraints that can be established. More efforts are required for a better understanding. Ideas are only briefly presented here in the hope of leading to other better suggestions from the community.
The general ideas are: i) if gravity is the only interaction between dark matter particles, the propagation range with a scale-independent rate of the cascade may extend to the smallest scale (free streaming scale in Fig. 13). If the free stream scale is comparable to the particle mass, the scaling laws we identified may provide useful insights into the properties of dark matter particles (Xu, 2022b); ii) the relevant physical quantities (waiting time, etc.) for the particle random walk and migration involves the particle mass $m_{p}$ (Eq. (72)). New constraints on particle mass can be identified if we can identify any constraints on these quantities. We will illustrate the second idea in this section:
@@
    The waiting time $\tau_{h}^{*}\equiv\tau_{hr}(r=r_{h}^{*})$ represents a time scale for a real physical process, i.e., the waiting time for the migration of particles on the halo boundary (Eq. (58)). It is reasonable to assume that the waiting time $\tau_{h}^{*}\geq t_{p}$, where $t_{p}=5.4\times 10^{-44}$s is the Planck time, the smallest possible unit of time for any physical process. With this and $\gamma_{r}\approx 0.5$ (Eq. (51)), the particle mass should satisfy
@@
    where $\dot{m_{h}^{*}}$ is the rate of mass accretion. On average, characteristic haloes accrete one particle of mass $m_{p}$ during an average waiting time $\tau_{g}^{*}$. For a typical mass $m_{h}^{*}(z=0)\approx 3.2\times 10^{13}M_{\odot}$ at the current epoch (Cooray and Sheth, 2002) and $t_{0}\approx 4.3\times 10^{17}$ s (the age of the universe), the rate of mass accretion $\dot{m_{h}^{*}}\approx 1.5\times 10^{26}$ kg/s. The particle mass $m_{p}$ can be related to the waiting time as
@@
    where $s_{h}^{*}$ is approximately the jumping length for the migration of particles on the halo boundary. Using Eq. (95), the speed of halo growth is related to that jumping length as
@@
    Especially we estimate the speed of halo growth and the jumping frequency at scale $r$ as
@@
    where $c=3\times 10^{8}$m/s is the speed of light. Plugging in numbers into Eq. (103) and considering that $\alpha\leq 1$ and $\beta\geq 1$ ($\dot{m_{h}^{*}}\approx 1.5\times 10^{26}$ kg/s, $\Omega_{DM}\approx$0.2449, and $Ht$=2/3), we should have
@@
All these constraints (Eqs. (90), Eq. (93), (100), and (107)) exclude the standard WIMPs and strongly suggest a heavy dark matter scenario, where superheavy right-handed neutrinos might be a good candidate. Detailed discussion on the possible properties of dark matter particles from the mass and energy cascade is presented in a different article (Xu, 2022b). Let us take the particle mass $m_{p}=10^{12}$eV in Eq.(107) as an example; all other quantities can be easily obtained from the scaling laws (Eq. (70)) and equations in this section. Table 2 summarizes these values of relevant physical quantities on the smallest and largest scale of $r$ in haloes of characteristic mass $m_{h}^{*}$ at the current epoch. For $m_{p}=10^{12}$eV, on the halo boundary, the waiting time $\tau_{gr}=\tau_{hr}=\Delta_{c}t_{p}$ and the jumping length $s_{h}^{*}=l_{p}$.
Table 2: Relevant quantities in haloes of characteristic mass $m_{h}^{*}$ at $z$=0. The lower part of the table contains quantities dependent on the particle mass $m_{p}$.
@@
Waiting time (s) $\tau_{hr}$ $\propto r^{-1}$ $1\times 10^{-6}$ $1.1\times 10^{-41}$

### v4

@@ 4.3 Particle random walk and energy cascade
Figure 13 presents a schematic diagram for the random walk of particles in haloes with a fluctuating gravitational potential, a mean field representation for halo structure formation (Section 4.1). Due to the discrete particle nature of haloes, the gravitational potential $\Phi$ in the haloes can never be smooth, as shown by N-body simulations. The total potential on scale $r$ can be decomposed into the mean and fluctuations. Here, $\Phi(r)=\Phi_{r}(r)+E_{r}(r)$, where $\Phi_{r}$ is the mean potential at scale $r$, and $E_{r}$ is the fluctuation of the potential at $r$. The mean potential, potential fluctuation, and density at scale $r$ read
$\begin{split}\Phi_{r}(r)=-\frac{Gm_{r}}{r},\quad E_{r}(r)=\frac{Gm_{p}}{d_{r}}\quad\text{and}\quad\rho_{r}(r)=\frac{m_{p}}{d_{r}^{3}},\end{split}$ (73)
where $m_{r}$ is the mass enclosed by $r$, $d_{r}$ is the mean spacing between particles on scale $r$, and $m_{p}$ is the mass of a single particle. The
Next, we focus on the average waiting time for the random walk of particles in a fluctuating potential, for example, jumping from position "1" to "2" in Fig. 13. The local temperature $T_{r}$ and the probability $P_{r}$ of a successful attempt read
[…]
where $\gamma_{r}$ is a numerical factor. The typical velocity $v_{r}$ on scale $r$ is related to the local potential $\Phi_{r}$ by the Virial theorem. Hereafter, we adopt the subscript ’r’ to represent quantities on scale $r$, the subscript ’p’ for these quantities on the smallest scale in haloes that is determined by the particle mass and dependent on the resolution of the N-body simulations, the subscript ’s’ for these quantities at scale radius $r_{s}$, the subscript ’h’ for these quantities on the scale of halo size $r_{h}$, and the superscript ’*’ for quantities in haloes of characteristic mass $m_{h}^{*}$.
Since potential fluctuations are usually much smaller than the mean potential ($m_{p}E_{r}\ll k_{B}T_{r}$), the probability of a successful attempt and the waiting time $\tau_{gr}$ for particle random walk read
where $\tau_{gr}(r)$ is the waiting time for the random walk at location $r$ that is inversely proportional to the local potential $\Phi(r)$, as we assumed in Section 4.1. Here, $\Gamma_{r}(r)$ is the attempt frequency at $r$ (similarly to the attempt frequency in solid diffusion), i.e., the number of attempts per unit of time for a particle to jump to a neighboring position. The waiting time $\tau_{gr}$ can be related to the waiting time $\tau_{g}$ for the random walk of haloes in mass space (Eq. (30)), i.e. $\tau_{g}\equiv\tau_{gh}=\tau_{gr}(r=r_{h})$.
These two equations lead to
where $m_{h}^{*}$ and $r_{h}^{*}$ are the mass and size of the haloes and $u_{h}^{*}=\Gamma_{r}d_{r}$ represents the speed of growth of the haloes. Here, we use the fact that $m_{h}^{*}\propto t$, $r_{h}^{*}\propto t$, and $\tau_{g}^{*}\propto t^{0}$ (Eq. (32)). For a typical characteristic mass $m_{h}^{*}\approx 3\times 10^{13}M_{\odot}$, the size $r_{h}^{*}\approx 1$Mpc at $z=0$, and the age of the universe of 13.7 billion years, the velocity $u_{h}^{*}$ is around 80km/s. Since the speed of halo growth $u_{h}^{*}$ remains constant while the density of characteristic haloes $\rho_{h}^{*}\propto a^{-3}$ decreases with cosmic time, we estimate that $u_{h}^{*}$ is independent of the halo density and should also be roughly independent of the scale $r$. The larger the mean particle distance $d_{r}$, the smaller the attempt frequency $\Gamma_{r}$ (also see Eq. (109)). Therefore, the waiting time of the particles is inversely proportional to the local potential, that is, $\tau_{gr}\propto\Phi_{r}^{-1}$ from Eq. (75). Similarly, the waiting time for haloes in mass space is inversely proportional to the halo potential, i.e., $\tau_{g}\propto\Phi_{h}^{-1}$ in Eq. (30).
Now, we focus on the change in particle energy associated with that elementary jump during infinitesimal time $dt_{r}$
$\begin{split}&\Delta E_{2}=\frac{1}{2}\left(1+\frac{2}{n_{e}}\right)m_{p}dv_{r}^{2},\end{split}$ (78)
where $n_{e}$ is the effective potential index ($n_{e}$=-10/7 on global scales in Section B). Here, $dv_{r}^{2}$ is the infinitesimal difference in the specific kinetic energy between two positions "1" and "2" in Fig. 13. We assume that the particle is in virial equilibrium before and after every jump. With $n_{e}>-2$, $\Delta E_{2}<0$ means that the particle energy decreases after this elementary jump. Similarly to the merging of haloes (Eq. (42)), the migration of particles also leads to the energy "dissipation". Both merging of haloes and migration of particles provide small-scale channels that contribute to the global-scale energy "dissipation" (Fig. 17). Now we introduce a key parameter $\varepsilon$ to represent the rate of change in the specific energy of the particle
$\begin{split}\varepsilon=\frac{\Delta{E_{2}}}{m_{p}dt_{r}}=\frac{1}{2}\left(1+\frac{2}{n_{e}}\right)\frac{dv_{r}^{2}}{dt_{r}},\end{split}$ (79)
where $\varepsilon<0$ represents the direction of energy transfer from small to large scales (inverse cascade). This is the microscopic origin of the definition of the parameter $\varepsilon$ in Eq. (70).
Next, we need to identify the microscopic counterpart of terms $dv_{r}^{2}$ and $dt_{r}$ in Eq. (79). Considering a thin spherical shell of thickness $s_{r}$ located at $r$ that contains $N_{r}$ particles. We will identify the waiting time $\tau_{hr}(r)$ for a successful jump for any one of these $N_{r}$ particles in that spherical shell. In analogy to Eq. (34) for halo random walk in mass space, we have $\tau_{gr}(r)=N_{r}\tau_{hr}(r)$ and the waiting time
where $s_{r}$ is the thickness of the spherical shell or the jump length of the random walk. Similarly to Eq. (35) for random walk in mass space, the corresponding rate of mass flow due to the outward migration of particles should read as follows:
i.e., a mass of $m_{p}$ is transferred to larger scales during the time $\tau_{hr}$ by migration. The waiting time $\tau_{hr}$ is the microscopical quantity equivalent to infinitesimal time $dt_{r}$, as well as the quantity $dv_{r}^{2}$
$\begin{split}dt_{r}=\tau_{hr}\quad\textrm{and}\quad{dv_{r}^{2}}=\frac{2}{5}{dm_{r}}\frac{v_{r}^{2}}{m_{r}}=\frac{2}{5}\frac{m_{p}v_{r}^{2}}{m_{r}}.\end{split}$ (82)
Here, the particle mass $m_{p}$ is the infinitesimal mass $dm_{r}$, and we use the scaling law $v_{r}^{2}\propto m_{r}^{2/5}$ in Eq. (71). Inserting $dt_{r}$ and $dv_{r}^{2}$ in Eq. (79), we can express the rate of the energy cascade $\varepsilon$ as:
$\begin{split}\varepsilon=\frac{1}{5}\left(1+\frac{2}{n_{e}}\right)\frac{m_{p}v_{r}^{2}}{m_{r}\tau_{hr}}=\frac{1}{5}\left(1+\frac{2}{n_{e}}\right)\frac{Gm_{p}}{r\tau_{hr}}.\end{split}$ (83)
With a particle on scale $r$ migrating from position "1" to "2" (Fig. 13) during the waiting time $\tau_{hr}$, the specific energy of $m_{p}v_{r}^{2}/m_{r}$ is transferred across scale $r$. Therefore, the parameter $\varepsilon$ represents the energy flux associated with the migration across scale $r$.
The mass flux is also related to the characteristic velocity $v_{r}$, the typical velocity at $r$ (Eq. (74)), so that the mass flux also reads
$\begin{split}\dot{m}_{r}=4\pi r^{2}\rho_{r}v_{r}.\end{split}$ (84)
By equating the two mass fluxes in Eqs. (81) and (84), the thickness $s_{r}$ and waiting time $\tau_{hr}$ read
Other relevant quantities include the mean particle distance $d_{r}$, the attempt frequency $\Gamma_{r}$, the waiting time $\tau_{gr}$ and $\tau_{hr}$, and the thickness of the shell $s_{r}$. These quantities depend on the particle mass $m_{p}$ and the growth speed of haloes of characteristic mass $m_{h}^{*}$, i.e., $u_{h}^{*}=\Gamma_{r}d_{r}\approx$80 km/s. From Eqs. (75) to (85), on scales $r\leq r_{s}$,
$\begin{split}&d_{r}=\beta_{r}^{-\frac{1}{3}}m_{p}^{\frac{1}{3}}\varepsilon^{-\frac{2}{9}}G^{\frac{1}{3}}r^{\frac{4}{9}}\textrm{,}\quad\Gamma_{r}=\beta_{r}^{\frac{1}{3}}u_{h}^{*}m_{p}^{-\frac{1}{3}}\varepsilon^{\frac{2}{9}}G^{-\frac{1}{3}}r^{-\frac{4}{9}},\\ &\tau_{gr}=\frac{m_{p}}{\alpha_{r}\gamma_{r}u_{h}^{*}}\varepsilon^{-\frac{2}{3}}Gr^{-\frac{2}{3}}\textrm{,}\quad\tau_{hr}=\frac{3}{5}\beta_{h}\gamma_{r}m_{p}\varepsilon^{-1}Gr^{-1},\\ &N_{r}=\frac{5{u_{h}^{*}}^{-1}}{3\alpha_{r}\beta_{h}\gamma_{r}^{2}}\varepsilon^{\frac{1}{3}}r^{\frac{1}{3}},\quad s_{r}=\frac{{u_{h}^{*}}^{-1}}{\alpha_{r}\beta_{h}^{1/3}}m_{p}\varepsilon^{-\frac{1}{3}}Gr^{-\frac{1}{3}}.\end{split}$ (86)
Since these quantities involve particle mass $m_{p}$, information on these quantities may provide useful insights into the particle mass and properties of dark matter (Section 5). Specifically, for haloes of characteristic mass $m_{h}^{*}$, the rate of energy cascade $\varepsilon=\varepsilon_{u}\approx$-$10^{-7}$m2/s3. On scale $r=r_{s}^{*}\propto a^{3/2}$, with subscript ’s’ standing for the scale radius, relevant quantities satisfy
$\begin{split}&\rho_{s}^{*}\propto a^{-2}\textrm{,}\quad d_{s}^{*}\propto a^{2/3}\textrm{,}\quad\Gamma_{s}^{*}\propto a^{-2/3},\\ &\tau_{gs}^{*}\propto a^{-1}\textrm{,}\quad\tau_{hs}^{*}\propto a^{-3/2}\textrm{,}\quad N_{s}^{*}\propto a^{1/2},\quad s_{s}^{*}\propto a^{-1/2}.\end{split}$ (87)
From Eqs. (79) and (83), the rate of energy cascade can be related to the energy "dissipated" due to the merging of haloes ($\Delta E_{1}$) and the random walk of particles in haloes ($\Delta E_{2}$):
$\begin{split}\varepsilon\propto\frac{\Delta E_{2}}{m_{p}\tau_{hr}}\propto\frac{\Delta E_{1}}{m_{h}\tau_{hr}(r=r_{h})}.\end{split}$ (88)
4.4 The scale-to-scale energy cascade in haloes
This section quantifies the energy cascade in haloes from N-body simulations. The same theory can also be applied to the energy flow in galaxy bulges and the associated scaling laws for bulge mass, size, dynamics, and the evolution of supermassive black holes (Xu:2024-Cosmic-quenching-and-scaling-laws). In analogy to the formulation of the energy cascade in halo mass space (Section 3.4), we examine the energy cascade in all haloes of a given mass $m_{h}$. We start by introducing cumulative functions along the radial direction $r$. The cumulative mass function $\Lambda^{h}_{m}(m_{h},r,z)$ (similar to $\Lambda_{m}$ in Eq. (37)) represents the total mass above $r$ that is averaged for all haloes of the same mass $m_{h}$
$\Lambda^{h}_{m}(m_{h},r,z)=\int_{r}^{\infty}\rho_{r}\left(m_{h},r^{\prime},z)\right)4\pi r^{\prime 2}dr^{\prime},$ (89)
where $\rho_{r}$ is the average mass density for all haloes of mass $m_{h}$.
Next, similarly to the formulation in Section 3.4, we also decompose the velocity of the halo particles $\boldsymbol{\mathrm{v}}_{p}$ into the mean halo velocity, $\boldsymbol{\mathrm{v}}_{h}=\langle\boldsymbol{\mathrm{v}}_{p}\rangle_{h}$, and the velocity fluctuation, $\boldsymbol{\mathrm{v}}_{p}^{{}^{\prime}}$, that is, $\boldsymbol{\mathrm{v}}_{p}=\boldsymbol{\mathrm{v}}_{h}+\boldsymbol{\mathrm{v}}_{p}^{\prime}$ (Eq. (47)). Here, $\boldsymbol{\mathrm{v}}_{h}$ represents the velocity of that halo, the average velocity of all particles in the same halo. Consequently, the total kinetic energy $K_{p}$ of a given halo particle can be divided into $K_{p}=K_{ph}+K_{pv}$. The virial kinetic energy, $K_{pv}={\boldsymbol{\mathrm{v}}_{p}^{\prime}}^{2}/2$, is the contribution from the velocity fluctuation due to the intra-halo interactions on small scales that are in the nonlinear regime. Only this part of the kinetic energy is relevant for the energy cascade in haloes. We introduce a cumulative function $\Lambda^{h}_{pv}$ for $K_{pv}$
Figure 14: The variation of the rate of energy cascade $\Pi^{h}_{pv}(m_{h},r,z)$ (Eq. (92)) with radial scale $r$ at different redshifts $z$ for haloes of characteristic mass $m^{*}_{h}(z)$ and $m_{h}=0.044m^{*}_{h}(z)$. A scale-independent (independent of $r$) constant rate of $\varepsilon\equiv\Pi^{h}_{pv}(r<r_{s})$ can be identified for an inverse cascade of virial kinetic energy $K_{pv}$} from small to larger scales. There also exists a simultaneous direct cascade of potential energy from large to small scales at a rate of $7/5\varepsilon$ (Xu:2022-Postulating-dark-matter-partic). For haloes with characteristic mass $m^{*}_{h}$, the rate of energy cascade $\varepsilon(m_{h}^{*},z)\equiv\varepsilon_{u}\approx-10^{-7}m^{2}/s^{3}$.
In a certain range of scales $r<r_{s}$ (inner haloes), the characteristic time on small scales is very small compared to the time on large scales. The small-scale motion does not feel the large-scale motion directly, except through the energy flux $\varepsilon$ across scales. When a statistical equilibrium is established, similar to the energy cascade in the halo mass space, we expect the rate of energy flow $\varepsilon(m_{h},z)$ to be independent of the scale $r$, i.e.
This can be tested by N-body simulations.
where $\varepsilon_{u}\equiv\varepsilon(m_{h}^{*},z)$ is the rate of energy flow in haloes of characteristic mass $m_{h}^{*}$. The dimensionless parameter $\nu$ is defined as $\nu=({m_{h}}/{m_{h}^{*}})^{2/3}$ (Xu:2023-Dark-matter-halo-mass-functions-and). The rate of the energy cascade varies with the mass and time as $\varepsilon\propto m_{h}^{2/3}a^{-1}$, while the haloes evolve towards the limiting equilibrium. We only have a vanishing energy cascade or $\varepsilon=0$ in fully virialized haloes with $a\rightarrow\infty$ (usually small and old haloes), a limiting state that can never be reached.
Finally, for haloes of any mass, the rate of energy cascade in that halo can be conveniently related to the energy "dissipated" by merging (Eq. (45)). For haloes of characteristic mass $m_{h}^{*}$ or any mass $m_{h}$, that rate is
$\varepsilon_{u}\propto\frac{\Delta E_{1}}{m_{h0}^{*}\tau_{g}^{*}}\propto\frac{Gm_{p}}{r_{h0}^{*}\tau_{g}^{*}}\propto\frac{Gm_{h}^{*}}{r_{h}^{*}t_{0}}\quad\textrm{and}\quad\varepsilon(m_{h},z)\propto\frac{Gm_{h}}{r_{h}t_{0}},$ (95)
where $t_{0}$ is the age of the Universe, $m_{h0}^{*}\equiv m_{h}^{*}(z=0)$ and $r_{h0}^{*}\equiv r_{h}^{*}(z=0)$. From Eq. (45), it can be directly related to the energy released $\Delta E_{1}$ during every merging. The continuous halo merging drives the system continuously, releasing energy and maximizing global entropy during the structure evolution.
@@ 5 Constraints on dark matter particle mass
Previous sections examined the cascade of mass and energy both across halo mass scales and within individual halos, and discussed how these cascades may depend on the mass and microphysical properties of dark matter particles. In this section, we outline a few illustrative constraints that can, in principle, be derived from these ideas. The discussion that follows is necessarily preliminary; further work is required to quantify these constraints and test their robustness. Therefore, these remarks are presented as suggestions intended to stimulate community follow-up studies.
Two general ideas motivate the constraints. First, if gravity is the only interaction among dark matter particles, for gravitationally dominant haloes of characteristic mass $m_{h}^{*}$ with universal spectrum and scaling laws, a scale–independent cascade rate may persist down to the smallest physical scales (the free‑streaming cutoff illustrated in Fig. 11). When the free‑streaming scale approaches the intrinsic length scale of the particle, the scaling relations derived above could provide direct insight into the properties of the particle (Xu:2022-Postulating-dark-matter-partic; Xu:2023-Universal-scaling-laws-and-density-slope). Second, several key physical quantities that govern the effective random walk and migration of particles (e.g., the waiting time) depend explicitly on the particle mass $m_{p}$ (Eq. (86)). If independent bounds can be placed on these quantities, they can be translated into novel constraints on $m_{p}$. In the remainder of this section, we illustrate the second approach with several examples.
@@
    The waiting time $\tau_{h}^{*}\equiv\tau_{hr}(r=r_{h}^{*})$ represents a characteristic timescale for a real physical process, i.e., the waiting time for the migration of particles across the halo boundary (Eq. (80)). It is reasonable to assume that the waiting time $\tau_{h}^{*}\geq t_{p}$, i.e. is greater than the Planck time $t_{p}=5.4\times 10^{-44}$s, the smallest possible unit of time for any physical process. With this and $\gamma_{r}\approx 0.5$ (Eq. (72)), the particle mass satisfies
@@
    where $\dot{m_{h}^{*}}$ is the rate of mass accretion. On average, characteristic haloes accrete one particle of mass $m_{p}$ during an average waiting time $\tau_{g}^{*}$. For a typical mass $m_{h}^{*}(z=0)\approx 3.2\times 10^{13}M_{\odot}$ at the current epoch (Cooray:2002-Halo-models-of-large-scale-str) and $t_{0}\approx 4.3\times 10^{17}$ s (the age of the universe), the rate of mass accretion $\dot{m_{h}^{*}}\approx 1.5\times 10^{26}$ kg/s. The particle mass $m_{p}$ can be related to the waiting time as
@@
    where $s_{h}^{*}$ is approximately the jumping length for the migration of particles on the halo boundary. Using Eq. (103), the speed of halo growth is related to the jumping length as
@@
    Especially, from Eq. (107), we estimate the speed of halo growth and the jumping frequency at scale $r$ as
@@
    where $c=3\times 10^{8}$m/s is the speed of light. Plugging in numbers into Eq. (111) and considering that $\alpha\leq 1$ and $\beta\geq 1$ ($\dot{m_{h}^{*}}\approx 1.5\times 10^{26}$ kg/s, $\Omega_{DM}\approx$0.2449, and HtHt=2/3), we should have
@@
All these constraints (Eqs. (98), Eq. (101), (108), and (115)) exclude the standard WIMPs and strongly suggest a heavy dark matter scenario, where superheavy right-handed neutrinos might be a good candidate. Detailed discussion on the possible properties of dark matter particles from the mass and energy cascade is presented in a different article (Xu:2022-Postulating-dark-matter-partic). Let us take the particle mass $m_{p}=10^{12}$GeV in Eq.(115) as an example; all other quantities can be easily obtained from the scaling laws (Eq. (71)) and equations in this section. Table 1 summarizes these values of relevant physical quantities on the smallest and largest scale of $r$ in haloes of characteristic mass $m_{h}^{*}$ at the current epoch. For $m_{p}=10^{12}$GeV, on the halo boundary, the waiting time $\tau_{gr}=\tau_{hr}=\Delta_{c}t_{p}$ and the jump length $s_{h}^{*}=l_{p}$.
Table 1: Relevant quantities in haloes of characteristic mass $m_{h}^{*}$ at zz=0. The lower table contains quantities dependent on the particle mass $m_{p}$.
@@
Waiting time (s) $\tau_{hr}$ $\propto r^{-1}$ $\times 10^{-6} 1.1$$1\times 10^{-41}$